Problem 4-3 Calculating Present Values [LO 2] For each of the following, compute the present value: (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) Present Value Years Interest Rate Future Value $ 11 5 % $ 18,228 3 10 42,417 15 13 799,382 20 12 656,816
Answer #1

To find the PV of a lump sum, we use:

PV = FV / (1 + r)t

PV = $18,228 / (1.05)11 = $10,657.53

PV = $42,417 / (1.10)3 = $31,868.52

PV = $799,382 / (1.13)15 = $127,813.79

PV = $656,816 / (1.12)20 = $68,089.99